Homework: 7.2 Score: 0 of 1 pt 7.2.35 14 of 14 (0 complete) HW Score: 0%, 0 of 14 pts Question Help caffeine content per 12-ounce bottle of cola is 35 milligrams You want to test this claim Duning your tests, you of thirty 12-ounce botties of cola has a mean caffeine content of 32.9 milligrams with a standard deviation of 6 9 milligrams Al a002, can A company that makes cola drinks states that the mean you reject the companys claim? Complete parts (a) through (e) ??, ?? p232 9 Ha ?-329 Ha u 32 5 Ha ?.35 Ha ?#329 Ha ?-35 Ha la#35 5
Answer #1

Given datas

the claim of the company that the caffeine present per 12 ounce bottel (population mean) \mu =35

no of sample taken (n)=30

caffeine present per 12 ounce bottel found in sample (Sample mean) \bar{X}=32.9

Sample standard daviation S=6.9

\alpha =0.02

So the claim is \mu =35

a)so the hypothesis will be

(f)

H_{o}\rightarrow \mu =35

H_{a}\rightarrow \mu \neq 35

now we have to check the claim so for that

Z score =    Z=\frac{\bar{X}-\mu }{S}=\frac{32.9-35}{6.9}=0.30

now we have \alpha =0.02

for \alpha =0.02 the Z criticle value

Z_{c}=\pm 2.33

since it is two tailed test so we have to check that valu Z score is within this criticle region or not

since -2.33<0.30<+2.33

so our test static lie within the criticle region in the bell shape curve so we can say that we dont have sufficient prove to reject our alternative hypothesis .

conclusion :--As we are not able reject our assumption i.e. \mu \neq 35

so we can say that we have enough evidence to reject the companies claim (i.e. null hypothesis ) that\mu =35